### v3

@@ B.1.2 Nanoparticle characterization
B.1.3 Protein Corona obtention
@@
CNPs from a 4 mg/mL stock solution and SiO2NPs from a 50 mg/mL stock solution were introduced into low protein-binding 1.5 mL microtubes and diluted to 2 mg/mL, respectively. We weighed 4 mg of Fib using a microbalance and diluted it in PBS to prepare the protein stock solution, which was later transferred into the NP solution to achieve a diluted concentration of 2 mg/mL. Consequently, the NP-protein corona complexes were formed.
@@
A total volume of 500 $\mathrm{\mu}$L from the final solution was transferred to a 1 mm cell to perform the UVRR. For the DCS, 100 $\mathrm{\mu}$L from the final sample were used at each run. For the DLS (Fig. S7 in ESI), the in-situ corona sample was diluted ten times to get an optimal attenuation. For the CD and UV absorption, the sample was diluted four times to avoid saturation of the signal.
B.2 UV Resonant Raman measurements
@@
UVRR spectra were collected by using the synchrotron-based UVRR set-up available at the BL10.2-IUVS beamline of Elettra Sincrotrone Trieste (Italy). [68] The excitation wavelength at 226 nm was obtained by setting the energy of the emission of the Synchrotron Radiation (SR) light by regulating the undulator gap aperture. The incoming SR light was further monochromatized through a Czerny-Turner monochromator (750 nm focal length Acton SP2750, Princeton Instruments, Acton, MA, USA) equipped with a holographic grating with 3600 grooves/mm. The Raman signal of the protein solutions was collected in back-scattered geometry and analyzed via a single-pass Czerny-Turner spectrometer (Trivista 557, Princeton Instruments, 750 mm of focal length) equipped with a holographic grating at 1800 g/mm. A thermoelectric cooled CCD camera optimized for the UV range was used to detect the signal at the output of the spectrometer. The resolution was set at about 1.6 $\mathrm{cm^{-1}}$/pixel. Cyclohexane (spectroscopic grade, Sigma Aldrich) was used for the calibration of the wavenumber scale. The final radiation power on the samples was about 20 $\mathrm{\mu}$W. To avoid photodegradation during the measurements, the samples were continuously oscillated with a 1 Hz frequency and a path length of 1 mm. For the collection of the temperature-dependent measurements (in the temperature range 22-86C with a step of 4 C) on protein solutions, a sample holder equipped with a thermal bath coupled to a resistive heating system was used to control the temperature with a stability of ±$\pm 0.1^{\circ}$. The solution of protein with and without NP was measured at a concentration of 2 mg/mL (the same weight/molar concentration was kept for protein and NP).
B.3 Circular dichroism and UV absorption measurements
@@
CD and UV absorption spectra were recorded using a Jasco J-810 polarimeter equipped with a plug-n-play single-cell Peltier with temperature control. We measured the samples of protein solution without and with NP in PBS buffer, pH=7.4, at a concentration of 0.25 mg/mL. The spectra were collected in the temperature range 22-87C with steps of 2C. The same weight/mol concentration of NP and protein was kept. All the solutions were freshly prepared in a rectangular quartz cell with a 1 mm path length. Each CD and UV spectrum was collected in the range from 260 to 190 nm with an increment of 1 nm, a 20 nm/min scan rate, and a 1 nm bandpass. Measurements were performed under a constant nitrogen flow, and the spectra were averaged from 4 scans. For each set of temperature-dependent measurements, the spectra of the solvent (PBS and NP solution in PBS at a concentration of 0.25 mg/mL) were subtracted from the corresponding sample spectra.
B.4 Data Curation
@@
The data obtained through different spectral techniques form a dataset containing UVRR, CD, and UV absorption spectra of Fib alone or in conjunction with CNP or SiO2NP for each temperature T, ranging from 22 to 86C, taken at temperature intervals of $\Delta T=4^{\circ}$. This results in an overall number of 17 spectra for each type of spectroscopic method. The UVRR spectra encompass silent regions where there is no useful information. Through spectral truncation [69], these regions are removed, retaining the following five intervals (see Section 2.1.2), which are reported in Table 1.
@@
Table 1: The UVRR spectral intervals in cm-1 unit under scrutiny together with the corresponding significant feature of interest and number of spectral components (or spectral intensities) $n_{c}$. Note that the interval $1660-1670$ cm-1 corresponds to a region of bulk Fib where a peak disappears altogether.
@@
$2800-3800$ $632$ OH Band
$1730-1800$ $41$ C=O
$1660-1670$ $6$ Peak
$1530-1580$ $29$ Trp
$1310-1390$ $45$ Trp
In CD and UV absorption spectra, we exclude spectral intervals below 210 nm because their signals are saturated due to the strong buffer absorption in such a wavenumber region. Consequently, after this spectral truncation, the available data ranges from 210 to 260 nm, corresponding to a total number of spectral components (or spectral intensities) denoted as $n_{c}=51$.
It is worth noting that the spectral truncation applied to CD spectra should address the challenges encountered when attempting to predict the secondary structure of Fib in the presence of CNP using the full spectral range (i.e., 190-260 nm) with the BeStSel software [70, 28]. Specifically, while the software reasonably predicts zero content of $\alpha$-helix structure of bulk Fib at high temperatures, it fails to do so for the protein’s spectra in the presence of CNPs (see Figure S3 in ESI). In this regard, we also found that when regression through Gaussian processes (GPs) [71, 72] is applied to spectra of Fib with CNPs, the performance deteriorates, indicating an increase of noise in the spectral data. This analysis is reported in Section 5 of ESI.
Finally, we merge the spectra for each system type—bulk Fib, Fib with CNPs, and Fib with SiO2NP—to create new composite spectra that contain significant spectral information as discussed above. These new spectra are formed by concatenating the UVRR intensities of five spectral intervals (as reported in Table 1), starting from the bottom in ascending order, along with the respective UV intensities, and then appending the CD intensities. As a result, the composite spectra under investigation consist of 856 dimensions.
B.5 Definitions of Metrics
Denoting two spectra by the sets $\left\{(\nu_{1},I_{1}),(\nu_{2},I_{2}),\cdots,(\nu_{N},I_{N})\right\}$, $\left\{(\nu_{1},Z_{1}),(\nu_{2},Z_{2}),\cdots,(\nu_{N},Z_{N})\right\}$, where $\nu_{i}$ is the wavenumber and $I_{i},Z_{i}$ the corresponding intensities, their similarity (or equivalently their dissimilarity) will be measured by computing a suitable distance between the vectors: $\mathbf{I}=\left(I_{1},I_{2},\cdots,I_{N}\right),\mathbf{Z}=\left(Z_{1},Z_{2},\cdots,Z_{N}\right)$, where $N$ corresponds to the number of the spectral observations. In Table 2 we report the mathematical formulae of Euclidean and Manhattan metrics, commonly used for assessing the similarity between the spectra.
@@ B.5 Definitions of Metrics
Next, the family of Wasserstein distances $W_{p}$ ($p=1,\cdots$) arises from the Kantorovich formulation of Monge’s mass transport problem [73, 74, 75]. The optimal transport distance $W_{p}$ is defined between probability distributions for a given metric. Thus, given two probability distributions $f\left(\mathbf{x}\right)$ and $g\left(\mathbf{y}\right)$ ($\mathbf{x},\mathbf{y}\in\mathbb{R}^{n}$), the Wasserstein distance yields the minimum work required to transform one distribution to the other. In the present work, we shall consider the one form of Wasserstein distance $W_{1}\left(f,g\right)$, that is, $p=1$, which is defined as the minimum of the following expectation value [76]:

### v5

@@ 2.2.1 Experimental details
Protein Corona obtention
@@
CNPs from a 4 mg/mL stock solution and SiO2NPs from a 50 mg/mL stock solution were introduced into low protein-binding 1.5 mL microtubes and diluted to 2 mg/mL, respectively. We weighed 4 mg of Fib using a microbalance and diluted it in PBS to prepare the protein stock solution, which was later transferred into the NP solution to achieve a diluted concentration of 2 mg/mL. Consequently, NP-protein corona complexes formed after 10 minutes of incubation.
@@
A total volume of $500\mathrm{\mu}$L from the final solution was transferred to a 1 mm cell to perform the UVRR. For the DCS, $100\mathrm{\mu}$L from the final sample was used at each run. For the DLS (Fig. S2 in ESI), the in-situ corona sample was diluted ten times to get an optimal attenuation. For the CD and UV absorption measurements, the sample was diluted four times to prevent signal saturation.
2.2.2 UV Resonant Raman measurements
@@
UVRR spectra were collected by using the synchrotron-based UVRR set-up available at the BL10.2-IUVS beamline of Elettra Sincrotrone Trieste (Italy) [46]. The excitation wavelength at 226 nm was obtained by setting the energy of the emission of the Synchrotron Radiation (SR) light by regulating the undulator gap aperture. The incoming SR light was further monochromatized using a Czerny-Turner monochromator (750 nm focal length, Acton SP2750, Princeton Instruments, Acton, MA, USA), equipped with a holographic grating featuring 3600 grooves/mm. The Raman signal of the protein solutions was collected in back-scattered geometry and analyzed using a single-pass Czerny-Turner spectrometer (Trivista 557, Princeton Instruments, with a focal length of 750 mm) equipped with a holographic grating at 1800 grooves/mm. A thermoelectrically cooled CCD camera optimized for the UV range was used to detect the signal at the spectrometer’s output. The resolution was set at about $1.6\mathrm{cm^{-1}}$/pixel. Cyclohexane (spectroscopic grade, Sigma-Aldrich) was used for calibrating the wavenumber scale. The final radiation power on the samples was about $20\mathrm{\mu}$W. To avoid photodegradation during the measurements, the samples were continuously oscillated with a 1 Hz frequency and a path length of 1 mm. For the collection of temperature-dependent measurements (in the temperature range of 22-86C with a step of 4C) on protein solutions, a sample holder equipped with a thermal bath coupled to a resistive heating system was used to control the temperature with a stability of $\pm 0.1^{\circ}$C. The solution of protein with and without NP was measured at a concentration of 2 mg/mL (the same weight/molar concentration was kept for protein and NP). Five individual frames were collected from each sample and averaged in order to improve the signal-to-noise ratio of each Raman spectrum. We checked that no inconsistencies, irreproducibility, or photodamaging effects were observed during data acquisition. The UVRR spectra were subtracted from an almost flat background (2nd-order polynomial baseline) and smoothed using a Savitzky-Golay smoothing filter with a 5-point window width. No normalization has been used in the representation of the experimental Raman profiles.
2.2.3 Circular dichroism and UV absorption measurements
@@
CD and UV absorption spectra were recorded using a Jasco J-810 polarimeter equipped with a plug-n-play single-cell Peltier with temperature control. We measured the protein solution samples, both with and without NP, in PBS buffer (pH 7.4) at a concentration of 0.25 mg/mL. The spectra were collected in the temperature range of 22-87C with a step of 2C. The same weight/mol concentration of NP and protein was kept. All the solutions were freshly prepared in a rectangular quartz cell with a 1 mm path length. Each CD and UV spectrum was collected in the range from 190 to 260 nm with an increment of 1 nm, a scan rate of 20 nm/min, and a 1 nm bandpass. Measurements were performed under a constant nitrogen flow, and the spectra were averaged from 4 scans. For each set of temperature-dependent measurements, the spectra of the solvent (PBS and NP solution in PBS at a concentration of 0.25 mg/mL) were subtracted from the corresponding sample spectra. The experimental UV absorption and CD profiles have been smoothed using a Savitzky-Golay smoothing filter with a 15-point window width. No normalization has been used in the representation of the experimental CD and UV absorption spectra.
2.2.4 Data Curation
@@
The data obtained through different spectral techniques form a dataset containing UVRR, CD, and UV absorption spectra of Fib alone or in conjunction with CNP or SiO2NP for each temperature T, ranging from 22 to 86C, taken at temperature intervals of $\Delta T=4^{\circ}$C. This results in an overall number of 17 spectra for each type of spectroscopic method. The UVRR spectra encompass silent regions where no useful information is available. Through spectral truncation [47], these regions are removed, retaining the following five intervals (see Section 3.1.2), which are reported in Table 1.
@@
Table 1: The UVRR spectral intervals in cm-1 unit under scrutiny together with the corresponding significant feature of interest and number of spectral components (or spectral intensities) $n_{c}$. Note that the interval 166016701660-1670 cm-1 corresponds to a region of bulk Fib where a peak disappears altogether.
@@
280038002800-3800 $632$ OH Band
173018001730-1800 $41$ C=O
166016701660-1670 66 Peak
153015801530-1580 $29$ Trp
131013901310-1390 $45$ Trp
In CD and UV absorption spectra, we exclude spectral intervals below 210 nm because their signals are saturated due to the strong absorption of the buffer in this wavenumber region. Consequently, after this spectral truncation, the available data ranges from 210 to 260 nm, corresponding to a total number of spectral components (or spectral intensities) denoted as $n_{c}=51$.
It is worth noting that the spectral truncation applied to CD spectra should address the challenges encountered when attempting to predict the secondary structure of Fib in the presence of CNP using the full spectral range (i.e., 190-260 nm) with the BeStSel software [48, 49]. Specifically, while the software reasonably predicts zero content of $\alpha$-helix structure in bulk Fib at high temperatures, it fails to do so for the protein’s spectra in the presence of CNPs (see Fig. S3 in ESI). In this regard, we also found that when regression is performed through Gaussian processes (GPs) [50, 51] on the spectra of Fib with CNPs, the performance deteriorates, indicating an increase in noise in the spectral data. This analysis is reported in Section 4 of ESI.
Finally, we merge the spectra for each system type —bulk Fib, Fib with CNPs, and Fib with SiO2NP— to create new composite spectra that contain significant spectral information, as discussed above. These new spectra are formed by concatenating the UVRR intensities of five spectral intervals (as reported in Table 1), starting from the bottom in ascending order, along with the respective UV intensities, and then appending the CD intensities. As a result, the composite spectra under investigation consist of 856 dimensions.
2.2.5 Definitions of Metrics
A central concept in our ML approach is the similarity among different spectra. Denoting two spectra by the sets $\left\{(\nu_{1},I_{1}),(\nu_{2},I_{2}),\cdots,(\nu_{N},I_{N})\right\}$, $\left\{(\nu_{1},Z_{1}),(\nu_{2},Z_{2}),\cdots,(\nu_{N},Z_{N})\right\}$, where $\nu_{i}$ is the wavenumber and $I_{i},Z_{i}$ the corresponding intensities, their similarity (or equivalently their dissimilarity) will be measured by computing a suitable distance between the vectors: $\mathbf{I}=\left(I_{1},I_{2},\cdots,I_{N}\right),\mathbf{Z}=\left(Z_{1},Z_{2},\cdots,Z_{N}\right)$, where $N$ corresponds to the number of the spectral observations. In Table 2, we report the mathematical formulae of Euclidean and Manhattan metrics, commonly used for assessing the similarity between spectra.
@@ 2.2.5 Definitions of Metrics
Next, the family of Wasserstein (or Earth Mover’s) distances $W_{p}$ ($p=1,\cdots$) arises from the Kantorovich formulation of Monge’s mass transport problem [52, 53, 54]. The optimal transport distance $W_{p}$ is defined between probability distributions for a given metric. Thus, given two probability distributions $f\left(\mathbf{x}\right)$ and $g\left(\mathbf{y}\right)$ ($\mathbf{x},\mathbf{y}\in\mathbb{R}^{n}$), the Wasserstein distance yields the minimum work required to transform one distribution to the other. In the present work, we shall consider the one form of Wasserstein distance $W_{1}\left(f,g\right)$, that is, $p=1$, which is defined as the minimum of the following expectation value [55]: